OpenStudy (anonymous):

The length of a rectangle is 2 feet less than three times the width.The area is 40 square feet .Find the length of the rectangle.

OpenStudy (dumbcow):

Area = Length*width 40 = L*W L = 3W-2 Use substitution -> 40 = W(3W-2) solve for W